John Doe Irgendwo 23, 1234 Sonstwo

Tel.: 1234  Fax.: 5678  E-Mail: John@doe.com

John Doe, Sonstwo

Anschrift abc 12345 Ort

24. 5. 2010

Sehr geehrte Damen und Herren, öäüÖÄÜß§@ Mit freundlichen Grüßen

John Doe